Spirialtex Inc., announced aggressive growth plans over the next three years. Below are the expected investments in fixed and working capital and corresponding pro forma income statements for each of the next 3 years.

Year 1 2 3
Tax Rate 40% 40% 40%
Capital Expenditures 20 30 40
Change in NWC 5 6 7
Shares Outstanding 6 6 6
Sales (000,000) 150 200 250
COGS 75 100 125
Sales & Administrative Expense 15 20 25
Depreciation & Amortization 50 51 52
Total Operating Costs 140 171 202
Operating Income (EBIT) 10 29 48
Earnings Before Taxes 6 24 42
Taxes 2.4 9.6 16.8
Net Income 3.6 14.4 25.2
EPS 0.6 2.4 4.2


Spirialtex has a corporate cost of capital of 12% and is expected to have a long run growth rate of 4%. What is the horizon value of the firm?

Group of answer choices

$439.40

$309.40

$211.50

$177.11

$125.00
Answer #1

Horizon value of firm will start at the end of year 3.

Free Cash flow from Firm(FCFF) = EBIT(1-Tax rate) + Depreciation- Change in Net Working Capital - Net Capital Spending

FCFF for year 3 = 48*(1-0.40) + 52 - 7 - 40

= $33.80

Calculating Horizon Value at the end of year 3:-

Horizon Value = \frac{FCFF_3(1+g)}{(WACC-g)}

where, FCFF3 = $33.80

g = Long-term growth rate = 4%

WACC = cost of capital = 12%

Horizon Value = \frac{33.80(1+0.04)}{(0.12-0.04)}

Horizon Value of Firm is $439.40

Option 1

If you need any clarification, you can ask in comments.

If you like my answer, then please up-vote as it will be motivating